### v1

# Integrated modeling of boron powder injection for real-time plasma-facing component conditioning

Abstract
An integrated modeling framework for investigating the application of solid boron (B) powder injection for real-time surface conditioning of plasma-facing components (PFCs) in tokamak environments is presented. Utilizing the DIII-D impurity powder dropper (IPD) setup, this study simulates B powder injection scenarios ranging from milligrams to tens of milligrams per second, corresponding to boron flux rates of $10^{20}-10^{21}$ B/s in standard L-mode conditions. The comprehensive modeling approach combines EMC3-EIRENE for simulating the plasma background and the Dust Injection Simulator (DIS) for the ablation and transport of the boron powder particles. The results show substantial transport of boron to the inboard lower divertor, predominantly influenced by the main ion plasma flow. The powder particle size (5-250 $\mu$m) was found to be insignificant for the scenario considered. The effects of erosion and redeposition were considered to reconcile the discrepancies with experimental observations, which saw substantial deposition on the outer plasma-facing components. For this purpose, the WallDYN3D code was updated to include boron sources within the plasma domain and integrated into the modeling framework. The mixed-material migration modeling shows evolving boron deposition patterns, suggesting the formation of mixed B-C layers or predominantly B coverage depending on the powder mass flow rate. While the modeling outcomes at lower B injection rates tend to align with DIII-D experimental observations, the prediction of near-pure boron layers at higher rates has yet to be experimentally verified in the carbon environment of the DIII-D tokamak. The extensive reach of boron layers found in the modeling suggests the need for modeling that encompasses the entire wall geometry for more accurate experimental correlations. This integrated approach sets a precedent for analyzing and applying real-time in-situ boron coating techniques in advanced tokamak scenarios, potentially extendable to ITER.
keywords:
boronization, boron powder injection, real-time coatings, plasma-facing components, wall conditioning, boron layers
journal: Nuclear Materials & Energy
1 Introduction
The technique of real-time wall conditioning and coating (RTWC) became recently subject of numerous studies in tokamaks and stellarators [1, 2, 3, 4, 5, 6, 7, 8, 9]. Traditional wall conditioning (boronization) techniques involve the injection of hazardous borane gases and interrupting plasma operation, falling short in the context of long-pulse steady-state reactor operation. Moreover, they often fail to condition the divertor plasma-facing components (PFCs) as efficiently as the main chamber walls. For ITER, the strategic decision to replace the beryllium wall with a full-tungsten (W) one necessitates adjustments to the baseline scenario and additional risk-mitigation strategies to sustain high-plasma performance, especially during the critical Q=10 phase. In this context, investigating alternative methods like boron material injection could be beneficial. The evolution of RTWC research has been greatly facilitated by the development of the impurity powder dropper (IPD) technology, designed for releasing single or multi-species impurity powders into the plasma in real-time [10]. The IPD can dispense a wide range of impurities in a non-gaseous state, including but not limited to lithium, boron, boron nitride, and silicon. Comprehending the transport and deposition processes of locally released impurities in the scrape-off-layer (SOL) holds key importance in developing this technology for ITER and next-generation fusion devices.
In this paper, a new modeling framework for injecting solid boron, the primary material used for real-time wall conditioning in current tokamaks, stellarators, and heliotrons, will be presented. It introduces a comprehensive, integrated approach to model the transport, erosion, and deposition of boron powder, illuminating the extent of coverage achieved on the main PFCs. This will be crucial for evaluating the suitability of this technique for next-generation fusion devices like ITER and Fusion Power Plants (FPPs), which require novel solutions to maintain wall integrity and sustain high-performance plasmas during extended operations [11, 12, 13].
The EMC3-EIRENE code [14, 15] effectively simulates full torus plasma, impurity transport, and plasma-material interactions. Previous studies using this code have identified non-axisymmetric effects from atomic impurity sources in axis-symmetric configurations, such as nitrogen gas seeding in Alcator C-Mod and boron injection in DIII-D [16, 17]. The Dust Injection Simulator (DIS) code [18] is used to model dust transport, similar to the DUSTT code, which has been applied to intrinsic dust in 2D plasma backgrounds [19, 20] and injected dust in the 3D heliotron LHD [21].
To decipher the mechanisms that lead to B wall coatings through B powder injection, it is crucial to correlate deposition with transport. For this purpose, the WallDYN3D code was employed, which has been recently extended to 3D configurations to enable studies in both tokamaks with non-axisymmetric wall geometries, RMPs, or other symmetry breaking such as local gas puffs and inherently three-dimensional configurations like stellarators [22, 23, 24].
Figure 1: (a) Top-down view and cross-sectional view on boron source distribution resulting from ablation of 5-250 $\mu$m size powder particles along there trajectories calculated with DIS. (c) Boron flux at peak location (inner divertor) for 5-250 $\mu$m size boron powder particles originating from B powder injection (zero impurity recycling).
The forthcoming sections of this paper briefly introduce the three models used (Sec. 2 and 3), discuss the transport effects, and resulting B distribution without impurity recycling (Sec. 4), proceed with the analysis based on mixed-material modeling (Sec. 3 and 6), and finally, present a summary and conclusions (Sec. 7).
2 Coupled dust and plasma impurity transport modeling with DIS and EMC3-EIRENE
For the present numerical study, the 3D plasma fluid and kinetic neutral edge transport (EMC3-EIRENE) have been coupled with the Dust Injection Simulator (DIS). EMC3 solves a set of reduced Braginskii fluid equations for particles, parallel momentum, and energies for electrons and ions [14]. EIRENE solves the kinetic transport equations for neutral atoms and molecules, including collisional processes [15]. EMC3-EIRENE can handle full-torus computational grids, including complex field and wall geometries.
The Dust Injection Simulator (DIS) model was designed to study the dynamics of dust particles within three-dimensional, time-evolving plasmas [18]. It employs a comprehensive approach, integrating equations that account for the motion, charge balance, temperature changes, and mass evolution of the dust particles. These equations allow for the simulation of dust transport and ablation processes, including sublimation.
The equation of motion for dust particles of mass $m_{d}$, $m_{d}\frac{dv}{dt}=R_{f}$, can be described as such:
$R_{f}=F_{drag}+F_{E}+F_{grav}+F_{centr.}$ (1)
where $R_{f}$ describes the force acting on the powder particles consisting of the drag force exerted by the plasma ($F_{drag}$), electrostatic force ($F_{E}$), gravitational force ($F_{grav}$), and centrifugal force ($F_{centr.}$). Presently, $F_{E}$ does not contribute to the dust transport since EMC3-EIRENE does not yet solve the electric field self-consistently. The deposition depths in the plasma of the injected materials are determined by material-specific properties such as vapor pressure, heat sublimation, emissivity, secondary electron emission, material density, and boiling and melting points. Note that the present version of the model assumes perfect spherical symmetry of particles, neglecting effects such as the rocket force [25].
The powder migration and ablation generate a 3D distribution of atomic boron sources that contribute to a total boron influx (B/s), corresponding to the mass flow injection rate (mg/s). Once the neutral impurities are released and ionized, their transport is modeled through the trace impurity fluid module, which solves the following equations:
$\nabla_{\parallel}\cdot(n_{z}V_{z\parallel})+\nabla_{\perp}\cdot(-D_{imp}\nabla_{\perp}n_{z})=S_{z}$ (2)
$0=-\frac{1}{n_{z}}\frac{dp_{z}}{ds}+ZeE_{\parallel}+m_{z}\frac{V_{\parallel}-V_{z\parallel}}{\tau_{zi}}$
$+0.71Z^{2}\frac{dT_{e}}{ds}+2.6Z^{2}\frac{dT_{i}}{ds}$ (3)
$=F_{P}+F_{E}+F_{fr}+F_{th,e}+F_{th,i}$
where $S_{z}$ denotes the impurity ionization source for charge state $Z$, and $D_{imp}$ the anomalous particle diffusivity. The collision time $\tau_{zi}$ between impurities and background main ions is calculated as $\tau_{zi}=\frac{1.47\cdot 10^{13}m_{z}(\nicefrac{{T_{i}^{3}}}{{m_{i}}})^{0.5}}{(1+\nicefrac{{m_{i}}}{{m_{z}}})n_{i}Z^{2}ln\Lambda}$ ($ln\Lambda\approx 15$)[26].
In many plasma edge scenarios, electrostatic and pressure forces are negligible [27], simplifying equation 4 to focus solely on friction and ion thermal forces, independent of charge state $Z$. At steady-state, this simplifies to a flow velocity balance [27, 17]:
$V_{z\parallel}\approx V_{\parallel}-\alpha\frac{dT_{i}}{ds}=V_{fr}-V_{i-th}$ (4)
The flows $V_{fr}$ and $V_{i-th}$ result from friction and ion thermal forces, respectively. A one-dimensional core transport model [14] simulates the ionization of impurities within the core plasma. Despite the axis-symmetric magnetic equilibrium, the trajectories of powder particles and the resulting impurity ion distributions are three-dimensional.
3 Coupling impurity transport with erosion and deposition dynamics in WallDYN3D
In the WallDYN3D model, the evolution of surface density for materials on the reactor walls, crucial for assessing material behavior under plasma exposure, is described by differential equations that account for erosion, deposition, and material mixing processes. For this study, the WallDYN3D model was extended to include impurity sources located within the plasma domain resulting from the B powder particle evaporation. The B atomic sources provided by the EMC3-EIRENE/DIS calculations have been included in the WallDYN3D model as a charge state $qi$ resolved B influx of component $ei$ at each wall location $wk$: $\Gamma_{ei,wk,qi}^{Inj}$.
The total impurity influx into the plasma of element $ei$ from wall $wk$ is given by the sum of reflected, eroded/sublimated, and injected flux, which is summarized as such:
$\Gamma_{ei,wj,qi}^{In}$ $=\sum_{wk}m_{ei,wk,wj,qi}(\Gamma_{ei,wk}^{Ero}+\Gamma_{ei,wk}^{Refl})+\Gamma_{ei,wj,qi}^{Inj}$ (5)
$\Gamma_{ei,wk}^{Refl}$ $=\sum_{qi}(1-R_{ei,wk,qi})\Gamma_{ei,wk,qi}^{In}$
$\Gamma_{ei,wk}^{Ero}$ $=\sum_{em}\sum_{qm}\Gamma_{em,wk,qm}^{In}Y_{ei,em,wk,qm}$
where $m_{ei,wk,wj,qi}$ describes the charge state resolved redistribution and migration matrix, and $R_{ei,wk,qi}$, and $Y_{ei,em,wk,qm}$, describe the reflection and erosion yields, respectively.
The governing equation for the areal density (atoms/m2) time evolution of wall surface element $\delta_{ei,wk}$ is defined as:
$\frac{d\delta_{ei,wk}}{dt}=\sum_{qi}\Gamma_{ei,wj,qi}^{In}-(\Gamma_{wk,qi}^{Ero}+\Gamma_{ei,wk}^{Refl})+\Gamma_{wk,qi}^{Bulk},$ (6)
where $\Gamma_{wk,qi}^{Bulk}$ represents a bulk material reservoir, that maintains the areal density of all wall elements.
Figure 2: (a, b) Distributions of neutral boron ($B^{0}$) and (c-g) ionized boron states ($B^{+1}$ to $B^{+5}$) at the B powder injection location in the torus. (h) Parallel flow velocity profile (Mach number) with blue and red indicating flows toward the inner and outer divertor targets, respectively.
[…]
4 Full-torus modeling of boron powder injection and transport with EMC3-EIRENE and DIS
A representative L-mode scenario for real-time wall conditioning in the DIII-D tokamak has been implemented using the EMC3-EIRENE code. This implementation utilizes a full-torus (360) grid based on the equilibrium #$\#179894$ at 3500 ms, specifically designed to fully capture the migration of boron powder particles and the resulting distributions of boron sources and fluxes. Previously, this scenario was used to model boron transport from a single atomic source [17]. In the current implementation, boron is introduced through powder injection, and the ablation and transport of this boron dust are simulated using the DIS code. A toroidal cross-section of the plasma grid domain for EMC3-EIRENE is shown in figure 1(a). The impurity powder source is situated at the tip of the upper small angle slot (SAS) divertor, where it is connected via a vertical tube to the powder dropper. The simulation accounts for a 5 cm diameter gap at this dropping location in the divertor geometry.
Figure 3: Time evolution of boron flux distribution simulated using WallDYN3D: (a) Initial deposition profile considering first-flight deposition only (excluding reflections, erosion, and redeposition effects); subsequent panels (b-f) illustrate changes at t=0, 1, 10, 20, and 400 seconds, incorporating reflections, erosion, and redeposition dynamics.
In the first step, the plasma and neutral transport have been calculated for the reference scenario without impurity injection for the entire torus. An upstream density of $n_{up}=1.5\cdot 10^{19}$ m-3 was set based on experimental measurements. The anomalous particle and heat cross-field transport in the plasma edge was prescribed by $D_{\perp}=0.33$ m2s-1 and $\chi_{\perp}=1.0$ m2s-1 for the purpose of this study based on reference values [17]. The cross-field diffusivity of main plasma and impurity ions is assumed to be equal for simplicity.
In the second step, the powder injection through the SAS port was modeled with DIS. A mass flow rate of $5$ mg/s was assumed, corresponding to an atomic rate of $2.8\times 10^{20}$ B/s. Within the model, the gravitational force accelerates the powder particles downwards into the plasma, reaching particulate velocities of 5 m/s at the injection point. Random angular distribution is assumed to take deviations from perfect vertical trajectories due to collisions within the drop tube into account, which can lead to trajectories that have components in the radial and toroidal directions, slightly deviating from an ideal vertical trajectory.
Figure 1(a,b) displays the boron source distributions along powder particle trajectories for sizes ranging from 5 to 250 $\mu$m, covering typical experimental particle diameters. The evaporation usually occurs when the powder particles reach the separatrix, where the electron temperature and associated sublimation heat flux $Q_{subl}$ increase, increasing the powder particle temperature to the sublimation point. For the mass loss rate follows: $\frac{dm_{d}}{dt}\propto Q_{subl}$. Larger particles, such as those measuring 150 and 250 $\mu$m, follow a predominantly vertical trajectory due to gravity and may cross the separatrix before fully ablating. Conversely, the trajectories of smaller particles (5 $\mu$m) are primarily influenced by the drag and centrifugal forces exerted by the toroidal plasma flow, which pushes these unconfined particles outward.
[…]
A notable increase in boron density is found on the high-field side of the plasma. This inboard peaking of the B density is attributed to the dominant parallel friction force terms associated with the parallel main ion flow, as delineated in equations 2 and 2 and illustrated as Mach flow profile in Figure 2. The parallel main ion flow in the powder deposition region shows a dominant component toward the inboard target (blue region), i.e., pushing the boron impurities effectively poloidally counter-clockwise. Thus, the inward directional bias of the flow culminates in heightened boron densities and fluxes impacting the plasma-facing components on the high-field side.
The boron incident flux distributions on the inner and outer divertor targets reveal a substantial asymmetry, with over $95\%$ of the total boron flux directed towards the inboard target according to the trace impurity fluid model. The toroidally averaged boron fluxes are shown for all particle sizes in Figure 1(c). The particle size does not significantly affect this inboard/outboard flux imbalance, with the frictional forces ($\propto V_{fr}\propto M_{||}$) of the main ion flow primarily governing the transport of boron in the plasma scenario considered.
Figure 4: Comparative analysis of boron migration: (a) fluid model, (b) kinetic model, and (c) numerical differences between fluid and kinetic models.
However, this boron distribution only represents direct deposition from the injection source without considering any subsequent removal or migration upon contact with the target surfaces, i.e., impurities do not recycle. As a result, deposition is narrowly confined to the far SOL, specifically between $Z=-115$ and $-95\,\text{cm}$ on the inboard divertor. Additionally, toroidal asymmetries in the boron deposition on the outer divertor arise from the localized nature of the boron source and the specific orientation of the flux tubes connecting the source region to the divertor targets, as noted in [17].
Notably, the substantial in/outboard imbalance in the B fluxes would put this deposition technique into question for low-Z wall coatings and wall conditioning. Contrary to this indication, experimental observations demonstrate substantial boron coatings even within areas that appear to be non-wetted based on the EMC3-EIRENE/DIS calculations [2, 29]). Therefore, the final step in this integrated modeling workflow must consider the effects of subsequent erosion and redeposition processes.
Figure 5: Evolution of WallDYN3D poloidal profiles at the DiMES toroidal location, showing (a) boron and (b) carbon fluxes at a rate of $10^{20}$ B/s, and (c) boron and (d) carbon fluxes at $10^{21}$ B/s for times of 0, 1, 10, 20, and 400 seconds. These profiles account for reflection, erosion, and redeposition. The poloidal position of DiMES is indicated by a dashed line.
5 3D boron erosion, migration deposition modeling with WallDYN3D
The integration of (re)erosion and redeposition processes was achieved by coupling WallDYN3D with the EMC3-EIRENE/DIS plasma solution. For this study, WallDYN3D was expanded to include boron sources (150 $\mu$m) within the plasma domain alongside traditional material targets, enabling the simulation of boron from powder particles.
The source rates were scaled to $10^{20}$ B/s and $10^{21}$ B/s to explore how experimental injection levels of a few milligrams per second up to 20 mg/s may affect the surface distribution and chemical composition. Carbon, the predominant intrinsic impurity in the DIII-D tokamak due to its graphite PFCs, was included in the simulation. The sputtering yields for C, both chemical and charge exchange, were set at 1%, with the physical sputtering yield determined by the energies imparted by the equation $3qT_{e}+2T_{i}$. The chemical and charge exchange sputtering for B were neglected. Multiple WallDYN3D time steps were carried out to model the evolution of the areal densities for B and C on the PFCs within the plasma domain. In each step, equations (5) and (6) are solved, resulting in effective recycling of the impurities.
A 3D visualization of the boron fluxes impacting the divertor targets at an IPD injection rate of $10^{20}$ B/s, as simulated using WallDYN3D, is shown in Figure 6. This figure initially depicts the boron deposition without reflections, erosion, or redeposition (Figure 6(a)). It then progresses through time steps from 0 to 400 seconds, incorporating these effects (Figure 6(b-f)). Notably, after reaching equilibrium after approximately 20 seconds, there is a significant alteration in the boron flux distribution. The outboard divertor target experiences an increased flux compared to the inboard target, as detailed in figures 6(e,f).
In WallDYN3D, impurity transport is treated kinetically by default [28]. A comparison between kinetic and fluid modeling was conducted to investigate potential differences in the transport between the high and low-field sides of relatively light B particles. In this case, boron is sourced from the high field side target. In figure 4, the B flux mappings are shown for the kinetic (Figure 4a) and fluid (Figure 4b) approaches. The boron fluxes are shown as charge state integrated re-distribution matrices, simplifying the three-dimensional migration data into a more comprehensible two-dimensional format through weighted averaging. The diagonal terms indicate the dominating local redeposition of B on the inner and outer divertor targets. The matrices also reveal an absence of transport from the inner to the outer target. However, values in the bottom right sections indicate additional transport from the outer to the inner target. This behavior is consistent in both the fluid and kinetic treatment qualitatively and quantitatively, as the difference in figure 6c shows. Therefore, the remaining differences in the initial deposition between the two models can be mainly attributed to differences in the impurity transport of boron ionized near the injection location.
Figure 6: (a) Boron surface concentrations after 400 iterations for moderate ($10^{20}$ B/s) and high ($10^{21}$ B/s) boron source strengths. (b) XPS measurement depth profile on DiMES sample from similar experiment showing a 40% boron concentration (40% carbon, not shown: 20% oxygen and other impurities) [2].
It is important to note that the simulation domain encompasses only a significant portion of the divertor targets within the DIII-D tokamak rather than the entire wall structure. Simulation results indicate that approximately 30% of the boron flux is not accounted for within this domain, suggesting that it would be deposited on wall elements outside the plasma simulation domain.
6 Impact of boron source strength on mixed-material coating composition
The dynamics of the B-C surface composition are affected by erosion and reflection properties on wall elements. One-dimensional boron and carbon flux profiles at the toroidal DiMES location (dashed vertical line, [30]) are shown in figure 5 for time steps of 0, 1, 10, 20, and 400 seconds, with boron influx rates of $10^{20}$ and $10^{21}$ B/s. The WallDYN3D calculations indicate that boron flux increases significantly due to reerosion, migration, and redeposition while carbon flux decreases. At the lower boron rate, boron flux to the high-field side target (inner peak) increases over time from $\approx 6\times 10^{19}$ to $1.2\times 10^{20}$ m-1s-1, and more dramatically on the outboard side (outer peak) from $\approx 3\times 10^{19}$ to $9\times 10^{20}$ m-1s-1. The carbon flux sees only a slight reduction. Conversely, at the higher boron rate ($10^{21}$ B/s), peak levels stabilize quickly, within 1 second at the inboard target, with carbon levels dropping by about 10
Adjusting the boron injection rate modifies the surface composition of divertor targets, as shown in Figure 6. With a boron influx rate of $10^{20}$ B/s, the B surface concentrations on the inner and outer targets reach approximately 35%, with boron area densities ranging from $0.3-1.4\times 10^{16}$ cm-2. Increasing the influx rate tenfold to $10^{21}$ B/s significantly enriches the boron on the target surfaces, potentially forming nearly pure boron layers, with area densities between $1.8$ and $3.6\times 10^{16}$ cm-2. This saturation phenomenon contrasts with lower influx rates, where the surface composition sustains a mixed B-C coating, approximating a B concentration closer to experimental findings (with B, C concentrations each being 40%). This effect can be attributed to the differential sputtering rates of B and C, which minimize the overall sputtering yields. This effect and associated boron-carbon compositions have been observed in previous DIII-D IPD experiments and traditional boronization procedures in tokamaks like TEXTOR and ASDEX-Upgrade[2, 31, 32]. For IPD experiments, however, achieving a balanced B:C concentration typically required B powder rates of 10-20 mg/s or $10^{21}$ B/s. Discussions in the literature, such as in [2], highlight a marginal dependence of surface composition on boron influx rates, underscoring the challenges of attaining pure boron layers in environments with significant carbon presence. The discrepancies between modeling and experiment may be partially due to the incomplete coverage of the in-vessel wall in the simulation domain. Carbon wall elements beyond the divertor might contribute to the total C influxes due to additional plasma-wall interactions (e.g., main wall CX-erosion) not considered in the present study.
7 Summary and conclusions
A new comprehensive modeling approach for real-time wall conditioning and in-situ boron powder coating is presented, utilizing three advanced models: EMC3-EIRENE, DIS, and WallDYN3D. These models collectively address 3D plasma transport, powder particle migration and ablation, and mixed-material surface evolution. Ablation modeling with EMC3-EIRENE/DIS shows that boron particles of up to 150 $\mu$m evaporate at the separatrix, where they reach sublimation temperature. The transport analysis shows that initial B fluxes are higher on the inboard divertor target due to the directional frictional forces from the main ion flow in the sublimation region. This effect closely aligns with findings from EMC3-EIRENE simulations using a single atomic B source in the same plasma scenario previously [17], and is consistent with beryllium transport modeling for JET and ITER using ERO2.0 [33]. However, first-flight deposition based on EMC3-EIRENE/DIS alone neglects impurity recycling, resulting in significant asymmetries in the divertor B fluxes concentrated in a narrow, far SOL region. The powder particle mass was not found to affect the distribution of B fluxes between the inboard and outboard divertor targets for particle sizes 5-250 $\mu$m in the present plasma scenario.
The modeling was extended by taking additional mixed-material sources and sinks, i.e., erosion and redeposition, into account using the WallDYN3D code. The code was extended to include external impurity sources from powder injection. Quasi-equilibrium solutions obtained with WallDYN3D show a toroidally uniform distribution with B fluxes on the outboard divertor target exceeding those on the inboard target. The approach also replicates B surface concentrations on the outer divertor close to experimental findings. However, the model underpredicts the B fluxes necessary to achieve this balance, likely partially due to a limited plasma grid domain that mainly covers the divertor region and neglects potential additional C sources from the remaining wall. Therefore, extending the simulation grid for the whole wall will be considered crucial for future work.
This integrated modeling approach provides a new framework for studying real-time wall conditioning and in-situ coating in advanced tokamak scenarios. The results suggest that poloidal and toroidal coverage of plasma-facing components with boron can be achieved through single-point B powder injection within the limitations of the model and its implementation. The insights garnered from this work will inform similar studies planned for ITER[13].
Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards DE-AC02-09CH11466 (PPPL), DE-FC02-04ER54698 (DIII-D), and DE-AC05-00OR22725 (ORNL). The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.
Disclaimer
This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.